Solve. 5x – 0.89 = 4(x – 0.71)
Answer by edjones(8007)

5x – 0.89 = 4(x – 0.71)
5x-.89=4x-2.84
5x-4x-.89+.89=4x-4x-2.84+.89
x=-1.95
.
Ed